OpenStudy (anonymous):

evaluate the function f(x)=40x square and g(x)=10x sq. (f+g)(10)=

OpenStudy (anonymous):

(40x^2)+(10x^2) = 50x^2 * 10 = 100x^2

OpenStudy (anonymous):

I got \[20\sqrt{10x}\]

OpenStudy (anonymous):

\[20\sqrt{10x}\]

OpenStudy (anonymous):

20sq.10x

OpenStudy (anonymous):

f+g*10 square of 40 is 4 and sq of 10 is 2

OpenStudy (anonymous):

*10